OpenStudy (mysticfury):

The sum of the measures of the interior angles of a polygon is 1980°. How many sides does the polygon have?

OpenStudy (mysticfury):

@jabez177 @Awolflover1 @TOMY12345 @Maddy1251 @retirEEd

OpenStudy (tomy12345):

The formula for interior angle sum is 180(n - 2). Backtrack it from that. 180(n - 2) = 1980. 1980 / 180 = 11 = n - 2. 11 + 2 = 13 = n. The polygon has 13 sides.

OpenStudy (mysticfury):

thanks

OpenStudy (tomy12345):

u wel